# AlGaN photonic device platform

### v1

@@ Abstract
In the rapidly evolving area of integrated photonics, there is a growing need for materials to support the fast development of advanced and more complex optical on-chip systems. Present photonic material platforms have significantly progressed over the past years; however, at the same time, they face critical challenges. We introduce a novel material for integrated photonics and investigate Aluminum Gallium Nitride (AlGaN) on Aluminum Nitride (AlN) templates as a platform for developing reconfigurable and on-chip nonlinear optical devices. AlGaN combines compatibility with standard fabrication technologies and high electro-optic modulation capabilities with reduced loss in a broad spectrum, making it a viable material for advanced photonic applications. We engineer and grow light-guiding photonic layers and fabricate waveguides, directional couplers, and other on-chip devices. Our AlGaN ring resonators demonstrate tunability and high Q-factors due to the low-loss waveguiding. The comprehensive platform paves the way for nonlinear photon-pair generation applications, on-chip nonlinear quantum frequency conversion, and fast electro-optic modulation for switching and routing.
@@ I Introduction
Advanced classical and quantum photonic applications, such as photonic neuromorphic computing [1], quantum sensing [2] and quantum networking [3, 4], rely on photonic integrated circuits that enable the compact, efficient and high-rate implementation of a variety of optical functionalities. In addition to functionalities adapted from stand-alone optical devices, integrated circuits provide access to mode-multiplexing or routing [5, 6] and efficient fiber coupling. A comprehensive photonic device platform [7], therefore, integrates various components, such as low-loss waveguides, efficient directional couplers, spectral filters, and tunable Mach-Zehnder interferometers (Fig. 1). Another versatile component is the tunable ring resonator, essential for spectral filtering, providing strong light-matter interaction, and enabling nonlinear conversion and photon-pair generation [8, 9, 10]. Finally, due to the large variety of classical and optical components available, the heterogeneous integration of components not fabricated directly into the platform material, such as detectors [11, 12] or quantum light sources [13], is required. A photonic platform material should exhibit low optical losses for efficient light transmission, fast electro-optic modulation capabilities for effectively manipulating optical modes, and significant optical nonlinearities to enable advanced functionalities.
@@
Currently, various material platforms are being explored, each presenting distinct strengths and weaknesses in meeting the comprehensive requirements of a photonic platform. For the ultraviolet and visible spectra, there are several more or less established candidates as an alternative to silicon, which is not transparent in this range. Lithium niobate on insulator is one of the leading platforms due to its outstanding electro-optic coefficient, high optical nonlinearity, and recent progress in technological development [14, 15, 16, 17]. However, it faces several challenges, including the need for non-standard etching processes to fabricate optical components. These processes are incompatible with the conventional manufacturing standards, potentially impeding its scalability [18]. Moreover, despite the advances in epitaxial techniques, the industry still predominantly relies on smart-cut lithium niobate derived from bulk crystals, which typically suffer from optical damage. The limited thickness range (300 nm) of smart-cut lithium niobate restricts some applications, such as nanoscale integrated optics [19]. Silicon nitride (SiN), another established material platform, offers low losses [20] and is compatible with silicon manufacturing standards. SiN-based active optical elements, however, such as tunable MZIs, typically rely on thermo-optic modulation due to their limited electro-optic functionality [21]. Thermo-optic control is comparably slow and generally limited to kHz modulation rates [22]. Another nitride-based material, aluminum nitride (AlN) on sapphire (AlNoS), has been explored due to its relatively high electro-optic constant, its substantial nonlinear coefficient, and its broad transparency, and moderate electro-optic coefficient [23, 24, 25, 26, 27]. AlN on sapphire, however, has yet other limitations. It is grown by sputtering or metallorganic chemical vapor deposition (MOCVD), typically on a sapphire substrate. Consequently, the AlN layer suffers from a high defect density at the sapphire interface [28, 29, 30], leading to relatively high optical scattering and absorption losses. Furthermore, commercially available AlN suffers from significant thickness variation, making it challenging for the scalable fabrication of photonic components.
@@
In this study, we introduce a platform for photonic integrated circuits based on aluminum gallium nitride (AlGaN), where the AlGaN layer acts as a guiding photonic layer that can be patterned into various components. This approach addresses challenges in AlN on sapphire platforms, such as crystal defects at the sapphire-AlN interface. We reduce optical losses and improve device performance by the AlGaN layer, shifting the optical mode away from this interface. The ability to control the ratio of aluminum to gallium in the ternary alloy AlGaN allows for tailored properties of the photonic layer, such as bandgap and refractive index engineering, for specific applications.
Moreover, AlGaN’s wide band gap enables efficient light transmission across a broad spectrum, from UV-C ( 250 nm) to long-wave infrared. Doping AlGaN for electrical conductivity combines optical and electronic functionalities, enabling, for example, current-induced refractive index change [31, 32].
@@
Our AlGaN platform benefits from the advanced state-of-the-art of GaN, which has, over the past decade, evolved into a foundational material for electronic and electro-optic devices, including power electronics and light generation [33, 34]. GaN LEDs and lasers are widely used for producing blue and green light [35, 36]. The epitaxial growth of GaN/AlGaN/InGaN heterostructures on sapphire templates, a well-established practice, opens up possibilities for integrating light sources and electronics on the photonic platform.
@@
The inherent properties of AlGaN, such as its relatively high electro-optic coefficient, make it an ideal candidate for rapid modulation and the development of reconfigurable optical devices. Its significant second-order optical nonlinearity paves the way for on-chip nonlinear photonic devices, including parametric oscillators and sum or difference frequency generators. The generation of correlated photon pairs in the ultraviolet (UV) and visible spectrum has been proposed using an AlGaN/AlN integrated photonics platform via spontaneous four-wave mixing (SFWM) in an AlGaN microring resonator [37]. AlGaN has also been explored as a potential platform for stimulated Brillouin scattering devices such as a racetrack Brillouin laser [38].
@@
Only very few studies have utilized AlGaN as a photonic material. Li et al. [39] fabricated large, multimode waveguides with an AlN/GaN multiple quantum well core and AlGaN claddings. Although this is a notable achievement, the large size and sub-optimal sidewall roughness do not enable application for single-mode photonics. In addition, the wavelength range is limited to around 350 nm and longer. Bruch et al. [40] fabricated ring resonators on AlGaN nanomembranes transferred onto SiO/Si substrates. While their work demonstrated the versatility of AlGaN, using nanomembranes introduces difficulties with handling and scalability.
@@
Here, we develop heterogeneously grown AlGaN into a photonic platform. To fabricate photonic devices, we grow AlGaN, with , directly on high-temperature annealed AlN-on-sapphire templates with a reduced dislocation density[41]. Into this heteromaterial, we fabricate the fundamental constituents of photonic integrated circuits with electron beam lithography: waveguides, Mach-Zehnder interferometers, directional couplers, ring resonators, and fiber couplers.
II Wafer Growth and Characterization
@@
The core ingredient of our novel integrated photonics platform is the heterostructured AlGaN material (Fig. 2(b)). The AlN/AlGaN layer stack was epitaxially grown on a c-plane-oriented sapphire substrate with an offcut of 0.25° towards an m-plane. First, a 350 nm thick AlN layer is deposited using epitaxial magnetron sputtering. The sputtered material’s threading dislocation density (TDD) was decreased through high-temperature annealing (HTA), adhering to the process described by Miyake et al. [42]. During HTA, a temperature of 1700°C is maintained for a relatively short duration of 1h to prevent the formation of aluminum oxynitride on the AlN surface [28]. Subsequently, the rough sputtered and annealed surface is smoothened by metal-organic vapor phase epitaxy (MOVPE) growth of 50 nm AlN in a step flow growth regime. The AlN layer, with a total thickness of 400 nm, exhibited a TDD of , as estimated from high-resolution X-ray diffraction (HRXRD) measurements of the symmetric 0002 and skew-symmetric 10-12 -rocking-curves [43]. Atomic force microscopy revealed a surface root mean square roughness of 0.09 nm over a area.
@@
Next, a 250 nm thick AlGaN waveguide layer is grown by MOVPE in a 6 x 2-inch close-coupled showerhead reactor on the AlN template, utilizing trimethylaluminum, triethylgallium, and ammonia as source materials, with hydrogen as the carrier gas, and a growth temperature of 1015°C. In-situ reflectometry measurements verify the attainment of the targeted layer thickness of 250 nm. After AlGaN growth, we measure the surface RMS roughness to be 0.5 nm. This smooth surface supports integrating photonic devices as introduced here and enables the hetero-integration of superconducting nanowire single-photon detectors (SNSPDs) [11] and quantum devices. The strain relaxation towards the underlying AlN buffer and the Al-mole fraction x of the layers is determined by HRXRD using reciprocal space maps (RSM) of the 0004 and 1124 reflections. Each RSM shows two sharp peaks (Fig. 2(a)), corresponding to the AlN template and AlGaN layers. By analyzing these data, an Al-mole fraction x of 0.69 and a relaxation of the in-plane compressive strain of only 3 can be determined for the AlGaN layer. This shows that the AlGaN grew almost pseudomorphically on the AlN template, and the generation of misfit dislocations for strain relief can thus be excluded.
III Optical Material Properties
@@
We used the prism coupling technique as described in [44, 45] to quantify the optical losses in the as-grown thin films. It utilizes a rutile prism to couple a tunable white laser (SuperK Fianium) into the films. The losses are quantified by monitoring the decay of the scattered light with a digital camera. We characterize a 400 nm AlNoS sample and an AlGaN/AlN/sapphire wafer. The AlNoS supports a single lateral TE mode at 630 nm, whereas the AlGaN wafer exhibits two distinct modes, TE and TE, at the same wavelength. By fine-tuning the coupling angle, we isolate and analyze the losses of these modes, as reported in Fig. 2(d).
@@
The TE mode, sustained by the AlGaN wafer, exhibits optical losses that decrease with increasing wavelength, reaching a minimum value of 0.37 cm above 600 nm. In contrast, the TE mode of the AlGaN and the TE mode of the AlNoS exhibited higher losses, with minimum values of approximately 1.36 and 1.0 cm, respectively, to rise again for wavelengths over 625 nm. It is worth noting that both of these modes are localized at the interface between two different materials, and as such, they are sensitive to the interface’s inherent roughness. Additional sources of loss could be due to defects and color centers in AlN and are presumably concentrated at the AlN-sapphire interface as shown by TEM measurements by [28, 29, 30].
@@
AlN has defects comprising a mix of impurities, vacancies, and lattice defects. Factors like threading dislocations influence the formation of these defects, although the exact conditions for their formation are yet not fully understood [46, 47, 48],
IV Device Fabrication
@@
In the fabrication process of waveguides and ring resonators, we initially deposit a 200 nm thick SiO layer using Plasma Enhanced Chemical Vapor Deposition (PECVD) to serve as a hard mask. We then coat a 10 nm thick Ti layer via e-beam evaporation to act as a charge dissipation layer, followed by a 400 nm ZEP 520A e-beam lithography resist. The process steps are illustrated in Fig. 3(a). After patterning the resist, we transfer the pattern to the hard mask using a two-step reactive ion etching process: we utilize SF plasma to etch Ti and CF plasma to etch SiO. In the final etching step, we process the AlGaN layer using BCl:Cl:He at a ratio of 10:50:10 sccm, under 600 W ICP power and 100 W RF power, at a pressure of 1 Pa. The AlGaN to SiO etch rate ratio is approximately 3/1. We implement a laser-cut notch at the chip’s edge to facilitate precise cleaving. Fig. 3(b) presents an SEM image of a cleaved facet, revealing the AlN and AlGaN layers. We measure the waveguides to have a wall angle of 75. Fig. 3(c) shows the simulated power profiles of the fundamental TE mode of the waveguides at wavelengths of 632 nm and 785 nm. At 632 nm, 95, and at 785 nm, 90 of the optical power resides above the AlN-sapphire interface, hence well removed from this lossy boundary.
V Tunable Microring Resonators

### v2

# AlGaN/AlN heterostructures: an emerging platform for nonlinear integrated photonics

@@ Abstract
In the rapidly evolving area of integrated photonics, there is a growing need for materials that satisfy the particular requirements of increasingly complex and specialized devices and applications. Present photonic material platforms have made significant progress over the past years; however, each platform still faces specific material and performance challenges. We introduce a novel material for integrated photonics: Aluminum Gallium Nitride (AlGaN) on Aluminum Nitride (AlN) as a platform for developing reconfigurable and nonlinear on-chip optical systems. AlGaN combines compatibility with standard semiconductor fabrication technologies, high electro-optic modulation capabilities, and large nonlinear coefficients while providing a broad and low-loss spectral transmission range, making it a viable material for advanced photonic applications. In this work, we design and grow AlGaN/AlN heterostructures and integrate fundamental photonic building blocks into these chips. In particular, we fabricate edge couplers, low-loss waveguides, directional couplers, and tunable high-quality factor ring resonators to enable nonlinear light-matter interaction and quantum functionality. The comprehensive platform we present in this work paves the way for nonlinear photon-pair generation applications, on-chip nonlinear quantum frequency conversion, and fast electro-optic modulation for switching and routing classical and quantum light fields.
@@ I Introduction
Advanced classical and quantum photonic applications, such as photonic neuromorphic computing [1], quantum sensing [2] and quantum networking [3, 4, 5, 6, 7] rely on photonic integrated circuits that enable the compact, efficient, and high-rate implementation of a variety of optical functionalities. In addition to functionalities adapted from stand-alone optical devices, integrated circuits provide access to mode-multiplexing or routing [8, 9] and efficient fiber coupling. A comprehensive photonic device platform [10], therefore, integrates various components, such as low-loss waveguides, efficient directional couplers, spectral filters, and tunable Mach-Zehnder interferometers (Fig. 1). A tunable ring resonator is another particularly versatile component, which can be employed to spectrally filter different modes, to enhance light-matter interaction, or to enable nonlinear frequency conversion and photon-pair generation.[11, 12, 13]. Additionally, given the large variety of electronic and optical components readily available on different material platforms, like detectors [14, 15] or quantum light sources [16, 17], heterogenous integration is a key approach for building photonic integrated chips with a high level of complexity. An ideal photonic platform should, therefore, facilitate such a method while also exhibiting properties such as low optical losses, fast electro-optic modulation, and significant optical nonlinearities to route and manipulate the propagating photons with high efficiency.
@@
Different material platforms are currently being studied, each with its own advantages and disadvantages in meeting the requirements of a versatile, scalable, and reliable photonic platform. For the ultraviolet and visible spectral range, there are several established alternatives to silicon, which is not transparent in this range. Lithium niobate on insulator (LNOI) is one of the leading platforms due to its outstanding electro-optic coefficient, high optical nonlinearity, and recent progress in technological development [18, 19, 20, 21]. However, LNOI faces several challenges. Non-standard etching processes are needed to fabricate optical components. These processes are incompatible with the conventional manufacturing standards, potentially hindering its scalability [22]. Moreover, despite the advances in epitaxial techniques, research, and industry still predominantly rely on smart-cut lithium niobate derived from bulk crystals, which typically suffer from crystal damage. Furthermore, the limited thickness range (300 nm) of smart-cut lithium niobate restricts some applications, such as nanoscale integrated optics [23].
Silicon nitride (SiN), another established material platform, offers low losses [24] and is compatible with silicon manufacturing standards. SiN-based active optical elements, however, such as tunable Mach-Zehnder Interferometers (MZIs), typically rely on thermo-optic modulation due to their limited electro-optic functionality [25]. Thermo-optic control is comparably slow and generally limited to kHz modulation rates [26]. Another nitride-based material, aluminum nitride (AlN) on sapphire, has been explored due to its substantial nonlinear coefficient, its broad transparency range, and its moderate electro-optic coefficient [27, 28, 29, 30, 31]. AlN on sapphire, however, has yet other limitations. It is usually grown by sputtering or metal-organic vapor phase epitaxy (MOVPE), typically on a sapphire substrate. Consequently, the AlN layer suffers from a high defect density at the sapphire interface [32, 33, 34], which leads to relatively high optical scattering and absorption losses. Furthermore, commercially available AlN suffers from non-negligible thickness variation, making it challenging to fabricate photonic components scalably.
@@
In this study, we introduce a platform for photonic integrated circuits based on aluminum gallium nitride (AlGaN), where the AlGaN layer acts as a guiding photonic layer that can be patterned into a large variety of optical components. This approach overcomes challenges in AlN on sapphire platforms, such as crystal defects at the sapphire-AlN interface and lateral thickness variation, and introduces further material-specific functionality. With the additional AlGaN layer, we reduce optical losses and improve device performance compared to AlN on sapphire by shifting the optical mode away from this interface. The ability to control the ratio of aluminum to gallium in the ternary alloy AlxGa1-xN allows for tailored properties of the photonic layer, such as bandgap and refractive index engineering. Moreover, AlGaN’s wide band gap enables efficient light transmission across a broad spectrum, depending on composition from below 250 nm to long-wave infrared. Doping AlGaN for electrical conductivity combines optical and electronic functionalities, enabling, for example, current-induced refractive index change [35, 36].
@@
Our AlGaN platform benefits from the advanced state-of-the-art of GaN, which has, over the past decade, evolved into a foundational material for electronic and electro-optic devices, including power electronics and light generation [37, 38]. For example, GaN LEDs and lasers are widely used for producing blue and green light [39, 40]. The epitaxial growth of GaN/AlGaN/InGaN heterostructures on sapphire templates, a well-established practice, opens up possibilities for integrating light sources and electronics on the photonic platform.
@@
The inherent properties of AlGaN, such as its relatively high electro-optic coefficient [41, 42], make it an ideal candidate for rapid modulation and the development of reconfigurable optical devices. Its significant second-order optical nonlinearity paves the way for on-chip nonlinear photonic devices, including parametric oscillators and sum or difference frequency generators. The generation of correlated photon pairs in the ultraviolet (UV) and visible spectrum has been proposed using an AlGaN/AlN integrated photonics platform via spontaneous four-wave mixing (SFWM) in an AlGaN microring resonator [43]. AlGaN has also been explored as a potential platform for stimulated Brillouin scattering devices such as a racetrack Brillouin laser [44].
@@
To the best of our knowledge, only very few studies have used AlGaN as a photonic material. Li et al. [45] fabricated large, multimode waveguides with an AlN/GaN multiple quantum well core and Al0.1Ga0.9N claddings. Although this is a notable achievement, the large size and relatively high sidewall roughness do not enable application for single-mode photonics. In addition, the wavelength range is limited to around 350 nm and longer. Bruch et al. [46] fabricated ring resonators on Al0.1Ga0.9N nanomembranes transferred onto SiO2/Si substrates. While their work demonstrated the versatility of AlGaN, using nanomembranes introduces difficulties with handling and scalability. Recently, Shin et al. showcased the enhancement of the Pockels effect in multiple AlGaN/AlN quantum wells on AlN due to the large built-in polarization of the quantum wells [47]. Despite encountering relatively high propagation losses (18.3 dB/cm), their work achieved a 20-fold increase in second-order susceptibility compared to bare AlN. This work shows the capability of the AlGaN/AlN as a platform for cutting-edge engineering to achieve nonlinearities beyond those available in bulk materials.
@@
Here, we develop heterogeneously grown AlGaN into a photonic platform. To fabricate photonic devices, we grow AlxGa1-xN, with x, directly on high-temperature annealed AlN-on-sapphire templates with a reduced dislocation density [48]. Into this heterostructure, we subsequently fabricate the fundamental constituents of photonic integrated circuits with electron beam lithography and plasma etching: waveguides, directional couplers, ring resonators, and fiber couplers. Finally, we present the design of a device to achieve efficient and flexible phase matching for entangled photon generation using Spontaneous Parametric Down-Conversion (SPDC) at telecom wavelengths employing such a heterostructure.
II Results
II.1 Wafer Growth and Characterization
@@
The core ingredient of our novel integrated photonics platform is an AlGaN heterostructure (Fig. 2(b)). The AlN/AlGaN layer stack is epitaxially grown on a c-plane-oriented sapphire substrate with an offcut of 0.25° towards an m-plane. First, a 350 nm thick AlN layer is deposited using epitaxial magnetron sputtering. The sputtered material’s threading dislocation density (TDD) was decreased through high-temperature annealing (HTA), following the process described by Miyake et al. [49]. During HTA, a temperature of 1700 °C is maintained for a relatively short duration of 1h to prevent the formation of aluminum oxynitride on the AlN surface [32]. Subsequently, the rough sputtered and annealed surface is smoothened by MOVPE growth of 50 nm AlN in a step flow growth regime. The AlN layer, with a total thickness of 400 nm, exhibits a TDD of , as estimated from high-resolution X-ray diffraction (HRXRD) measurements of the symmetric 0002 and skew-symmetric 10-12 -rocking-curves [50]. Atomic force microscopy (AFM) revealed a surface RMS roughness of 0.09 nm over a area.
@@
Next, a 250 nm thick AlGaN waveguide layer is grown by MOVPE in a 6 x 2-inch close-coupled showerhead reactor on the AlN template, utilizing trimethylaluminum, triethylgallium, and ammonia as source materials, with hydrogen as the carrier gas, and a growth temperature of 1015 °C. In-situ reflectometry measurements verify the attainment of the targeted layer thickness of 250 nm. After AlGaN growth, we measure the surface RMS roughness by AFM to be 0.5 nm. This smooth surface supports integrating photonic devices as introduced here and enables the hetero-integration of superconducting nanowire single-photon detectors (SNSPDs) [14] and quantum devices. The strain relaxation towards the underlying AlN buffer and the Al-mole fraction x of the layers is determined by HRXRD using reciprocal space maps (RSM) of the 11-24 reflection. Each RSM shows two sharp peaks (Fig. 2(a)), corresponding to the AlN template and AlGaN layers. By analyzing these data, an Al-mole fraction x of 0.69 and a relaxation of the in-plane compressive strain of only 3 can be determined for the AlGaN layer. This shows that the AlGaN grew almost pseudomorphically on the AlN template, and the generation of misfit dislocations for strain relief can thus be largely excluded.
II.2 Optical Material Properties
@@
We use the prism coupling technique as described in [51, 52] to quantify the optical losses in the as-grown thin films. It utilizes a rutile prism to couple a tunable white laser (SuperK Fianium) into the heterostructures. The losses are quantified by monitoring the decay of the scattered light with a digital camera. We characterize and compare a 400 nm AlNoS sample and an AlGaN/AlN/sapphire wafer. The AlNoS supports a single TE0 mode at 630 nm, whereas the AlGaN structure exhibits two distinct modes, TE0 and TE1, at the same wavelength. By fine-tuning the coupling angle, we isolate and analyze the losses of these modes, as reported in Fig. 2(d).
@@
The TE0 mode, in the AlGaN structure, exhibits optical losses that decrease with increasing wavelength, reaching a minimum value of 2.3 dB/cm above 600 nm. In contrast, the TE1 mode of the AlGaN and the TE0 mode of the AlNoS exhibited higher losses, with minimum values of approximately 5.9 and 4.3 dB/cm, respectively, which increase again for wavelengths over 625 nm. It is worth noting that both of these modes are localized at the interface between two different materials and, thus, are sensitive to the interface’s inherent properties. Besides the roughness at the interface, additional sources of loss could be due to defects and color centers in AlN, which are presumably concentrated at the AlN-sapphire interface as shown by TEM measurements by [32, 33, 34].
@@
AlN has defects comprising a mix of impurities, vacancies, and lattice defects. Factors like threading dislocations influence the formation of these defects, although the exact conditions for their formation are yet not fully understood [53, 54, 55].
II.3 Device Fabrication
@@
In the fabrication process of waveguides and ring resonators, we initially deposit a 200 nm thick SiO2 layer using Plasma Enhanced Chemical Vapor Deposition (PECVD) to serve as a hard mask. We then coat a 10 nm thick Ti layer via e-beam evaporation to act as a charge dissipation layer, followed by 400 nm ZEP 520A e-beam lithography resist. The process steps are illustrated in Fig. 3(a). After patterning the resist, we transfer the pattern to the hard mask using a two-step reactive ion etching (RIE) process: we utilize SF6 plasma to etch Ti and CF4 plasma to etch SiO2. In the final etching step, we process the AlGaN layer using BCl3:Cl2:He at a ratio of 10:50:10 sccm, under 600 W ICP power and 100 W RF power, at a pressure of 1 Pa. The AlGaN-to-SiO2 etch rate ratio is approximately 3/1. We implement a laser-cut notch at the chip’s edge to facilitate precise cleaving. Fig. 3(b) presents an SEM image of a cleaved facet, revealing the AlN and AlGaN layers. We measure the waveguides to have a wall angle of 75∘. Fig. 3(c) shows the simulated power profiles of the fundamental TE mode of the waveguides at wavelengths of 632 nm and 785 nm. At the measurement wavelength of 785 nm used for our ring resonators, 90 of the optical power is located above the AlN-sapphire interface. At a shorter wavelength of 632 nm, this number increases to 95, indicating that most of the optical power is well separated from the lossy boundary at the sapphire interface.
II.4 Tunable Microring Resonators